OpenStudy (anonymous):

Find the solution(s) of x^2+3x-18=0

OpenStudy (anonymous):

\[(x+6)(x-3) = 0\]

OpenStudy (muzzack):

Two solutions were found : x = 3 x = -6

OpenStudy (skullpatrol):

Any questions?

OpenStudy (anonymous):

How do you find the solutions?

OpenStudy (mathmale):

cormacpayne and muzzack: good contributions. But would you kindly help the questioner find his or her own answers and do NOT find (let alone post) them for him/her yourself. We (or at least others on OpenStudy and I) want each experience a student has here on OpenStudy to be practice towards developing academic independence. Thank you.

OpenStudy (skullpatrol):

You find the solutions by solving the equation.

OpenStudy (anonymous):

How do I solve the equation?

OpenStudy (mathmale):

Quinn33: You have various methods available to you to solve this quadratic equation. cormacpayne applied one of the more common ones: he / she factored the quadratic. Other methods include the quadratic formula, graphing, completing the square, and so on. Could you copy down the original quadratic equation and factor it on your own, without help? If not, I'd suggest you ask questions and keep trying until you're able to do that.

OpenStudy (skullpatrol):

You have to find the values of the variable that will make the equation a true statement.

OpenStudy (anonymous):

I'm so lost.

OpenStudy (skullpatrol):

Ok, let me explain. What is zero times any number equal?

OpenStudy (anonymous):

0?

OpenStudy (skullpatrol):

Yes you are right.

OpenStudy (skullpatrol):

What value of x makes x-3 =0?

OpenStudy (anonymous):

0?

OpenStudy (skullpatrol):

Nope. Solve for x in x-3=0. What do you get?

OpenStudy (anonymous):

3?

OpenStudy (goformit100):

A Warm Welcome to Open Study. Stay Studious.

OpenStudy (skullpatrol):

Correct. So that means if you replace x with 3 you will get the true statement 3 - 3 = 0 right?

OpenStudy (anonymous):

right

OpenStudy (skullpatrol):

So what are the solutions to (x - 3)(x + 6) = 0?

OpenStudy (skullpatrol):

And the same idea holds for x + 6 = 0

OpenStudy (anonymous):

3 and -6?

OpenStudy (skullpatrol):

Yes, can you tell me why?

OpenStudy (anonymous):

Not really.

OpenStudy (skullpatrol):

What happens to (x - 3)(x + 6) = 0 when you replace x with 3?

OpenStudy (skullpatrol):

Try it and find out :)

OpenStudy (skullpatrol):

(x - 3)(x + 6) = 0 Let x =3 (3 - 3)(3 + 6) = 0 0(9) = 0 <----is this true?

OpenStudy (muzzack):

yes it is true 0=0

OpenStudy (skullpatrol):

Similarly, let x = -6 gives (x - 3)(x + 6) = 0 (-6 - 3)(-6 + 6) = 0 -9(0) = 0 <---is this true?

OpenStudy (mathmale):

Seeing that 0 = 0 in this case, to what conclusion can you come about x=-6 as a solution or not a solution of the given quadratic equation?

OpenStudy (skullpatrol):

We know x = 3 or x = -6 are solutions of (x - 3)(x + 6) = 0 BECAUSE zero times any number is zero. They are the only values of x that will make (x - 3)(x + 6) = 0 into a TRUE statement.

OpenStudy (mathmale):

I'd like to recognize skullpatrol for doing such a good job of guiding Quinn towards finding his/her own answers, and for his excellent interpretation of those results at the end. Way to go, skullpatrol!

OpenStudy (skullpatrol):

Thanks @mathmale :)